### v2

@@ 3.1 NTN PLATFORM SWAP LIMITATION
The typical weights of macro massive MIMO RU and baseband unit (BBU) are over 10Kg and 5Kg, respectively [Ericsson_MIMO]. Noticeably, in the 5G era, the maximum power consumption of a 64T64R active antenna unit (AAU) is 1000–1400 W, and that of a BBU is about 2000 W [Huawei_5Gpower]. However, different NTN platforms have various SWaP limitations, which should be considered in the NTN development stage [Haq_2025, shang2025novel].
@@
The DJI flagship untethered UAV platform Matrice 350 RTK achieves a loading capability of 2.7 Kg with 31 minutes of flight time, which cannot carry the macro BS equipment discussed above [DJI_350]. Although DJI has customized a delivery UAV platform with the 40 Kg loading capability (i.e., DJI Flycart30). The flight time is limited to 8 minutes, which fails to guarantee continuous connections to the users [DJI_flycart30]. Apart from that, the battery is originally designed to support the UAV flight with a typical capacity of 0.3 KWh (e.g., TB65 from DJI). Apart from untethered UAVs, tethered UAVs have also been regarded as an emerging NTN platform due to their tethered systems. The DG-M20 tethered UAV, from Dagong Technology Co., Ltd, can carry 15 Kg payloads hovering at 100 m for 24 hours [DG_tetheredUAV]. Furthermore, the HAP has also attracted attention from both academia and industry due to its relatively lower altitude and quasi-stationary position. Taking the HAP from Stratospheric Platforms Ltd as an example, with a wingspan of 60m, it can support a maximum of 140 Kg payload and can provide a 20 kW power supply [Strato_HAP].
@@
Considering the SWaP limitations above, the standardized function split and open interface in ORAN provide flexibility for the NTN MNO to customize the onboard components from different vendors, including O-RU, O-DU, and O-CU, to meet the NTN platform SWaP requirements. For example, the commercial untethered UAV can be utilized to implement a micro or pico network with integrated O-RU, O-DU, O-CU, and 5GC on board, which acts as an end-to-end flexible deployment option and avoids adding peripheral equipment for FH, midhaul, and backhaul transmissions. On the contrary, the power cable and optical fiber in the tethered UAV system allow it to serve as a macro network with various onboard options, including O-RU, O-RU+O-DU, and O-RU+O-DU+O-CU, which can be configured according to the latency requirement of the service.
3.2 NTN ROBUST CONNECTIVITY
@@
Due to the mobility of the NTN platforms and the varying channel conditions, one important challenge is to guarantee robust connectivity during operation, especially for the LEO satellites with high speeds. To solve this challenge, a simple solution is to deploy a massive constellation of satellites, which can provide redundant access links and transportation links under a highly dynamic network topology. For example, SpaceX from the USA has successfully launched more than 5000 Starlink satellites into orbit. Apart from that, China has proposed GW-A59 and GW-2 sub-constellations as GuoWang (national network) mega-constellation, a large satellite internet project aimed at providing broadband services. It is noted that GW-A59, with 6,080 satellites, is intended to orbit at a lower altitude (below 500 km), while GW-2, with 6,912 satellites, will orbit at a higher altitude (1,145 km above Earth). However, the massive deployment of satellites imposes a heavy financial investment burden on a single NTN MNO and requires long development periods.
@@
The integration of ORAN’s interoperability into NTN has the potential to bring new opportunities to solve the challenge above by enabling network sharing among MNOs. At the end of 2024, 3GPP has identified Indirect Network Sharing (INS) as the next evolution step towards 5G-Advanced Rel-20, which aims to enhance network co-construction and sharing capability [3gpp23501]. It is noted that 3GPP proposed INS as an effective solution to mitigate the compatibility issues between the network elements of sharing parties, e.g., shared satellite access network and two or more core networks of terrestrial participating operators via multiple operator core network (MOCN) [3gpp32130]. However, to avoid direct connections between network elements of different MNOs, INS is based on the build of a hosting network in advance by all participating MNOs, which cannot adapt to the high mobility and complex NTN environments. Therefore, ORAN surpasses the newly proposed INS concept by enabling direct connections among MNOs, and has the potential to achieve real-time on-demand sharing based on the instantaneous network status.
3.3 NTN PLATFORM CONTROL AND DATA ANALYTICS
@@
In traditional TNs, the BS is usually installed on fixed terrestrial infrastructure with a sufficient and stable power supply, such as cell towers, where the network performance is independent of the terrestrial infrastructure during operation. However, the instantaneous position, speed, attitude, and battery status of the NTN platforms impact the network performance significantly via handover, beam management, power allocation, and so on. Therefore, it is of vital importance to design an interface between the NTN platform and onboard communication components, which aims to achieve dynamic communication algorithm optimization, adapting to the instantaneous NTN platform status, and also achieve proactively NTN platform control for network performance enhancements. However, following the vendor-locked principle in traditional RAN, the vendor is required to develop both the NTN onboard components and the NTN platform to enable them to be compatible with each other, which will be a new challenge for development and financial investment.
@@
Although the latest O-RAN Alliance white paper has proposed to integrate ORAN into NTN to alleviate the challenge above, where ORAN is expected to provide a standardized and open interface from RIC/SMO to the NTN platform data [ORAN_white_paper]. The white paper mainly focuses on offloading the additional terrestrial traffic instead of designing the open interface between the onboard component and the NTN platform for real-time control and optimization. Therefore, it is important to further design open interfaces from the NTN platform to O-RU, O-DU, and O-CU, providing network performance enhancements and deployment flexibility. For example, DJI has provided standard application programming interface (API) functions to control UAV mobility and access real-time UAV status. Specifically, the Roll, Pitch, and Yaw information and battery status can be accessed periodically under pre-defined binary control and command (C&C) formats. Different from the ORAN in TN, NTN platform manufacturing communities are required to jointly design the above open interface with the O-RAN Alliance, hence to capture the characteristics of various NTN platforms.
3.4 NTN ON-ORBIT TEST AND UPGRADE
@@
Traditional RAN is based on vendor-locked hardware, and most of the network functions and algorithms are implemented in specialized hardware with limited upgrade capability, including application specific integrated circuit (ASIC), field-programmable gate array (FPGA), or digital signal processors (DSP). It is noted that specialized hardware works well on the terrestrial networks due to easy accessibility and relatively low costs for testing and upgrading. However, launching NTN for the testing or deployment of new features is much costly and time-consuming [sandri2023ns, liu2024challenges]. Therefore, it is important to enable the on-orbit test and upgrade in NTN, which brings flexibility and shortens the DevOps periods. It is noted that, based on network function virtualization (NFV), ORAN enables the algorithms to be deployed on commercial off-the-shelf (COTS) hardware, e.g., a central processing unit (CPU) and a graphics processing unit (GPU). Enabled by the NFV, the MNO can hence test and upgrade the NTN in an agile DevOps manner.
@@ 4 AI-NATIVE ORAN-BASED NTN FRAMEWORK
Motivated by the advantages of ORAN-based NTN and the envisioned AI-based 6G networks, in this section, we propose an AI-Native ORAN-based NTN framework, where the key technical challenges and potential solutions are presented.
Figure 6: Proposed orchestrated AI-Native ORAN-based NTN architecture.
4.1 ARCHITECTURE
@@
As illustrated in Fig. 6, the proposed framework spans the RAN, transport network (TPN), and core network (CN) domains. Specifically, the RAN domain is implemented within the NTN segment, while the TPN domain may include both NTN and terrestrial segments, and the CN domain is deployed in the terrestrial segment. Each domain is managed and orchestrated by an AI-based controller. More importantly, a cross-domain AI controller is introduced above the three single-domain controllers to enable end-to-end AI-driven optimization. It is worth noting that all controllers are equipped with digital twin technology to enhance data analysis and predictive capabilities. Accordingly, the following subsection analyzes the key technological enablers that facilitate the implementation of this framework.
4.2 KEY ENABLERS
4.2.1 DYNAMIC WIRELESS FH SPLIT
@@
The gNB’s operation is conceptually structured as a series of functions, with the distribution of these functions defined by functional splits. Within the 3GPP framework, eight distinct functional split options exist, as shown in Fig. 7. Each split option describes how the logical nodes interrelate to one another and what specific activities each undertakes [openRAN-8]. From Option 1 to 8, there is a higher degree of centralization and a notable improvement in resource management efficiency, which leads to reduced O-RU complexity and associated costs. However, this comes at the expense of imposing high data rates and strict latency requirements on the FH between O-DU and O-RU [larsen2018survey].
@@
Typically, the O-RAN Alliance lends its support to the Option 7.2x variant for networks with high-capacity and high-reliability requirements, further splitting the functions in the L-PHY based on the PHY split in Option 7. This option is made possible by using the eCPRI interface for FH. In this split, the O-CU/O-DU manages RRC, PDCP, RLC, MAC, H-PHY, and parts of L-PHY, while the O-RU handles RF and the rest of L-PHY. The ’x’ in this split name can be either ’a’ or ’b’, depending on whether some of the PHY functionalities reside in H-PHY or L-PHY [openRAN-7]. Without loss of generality, our following discussion is based on the downlink transmission function split. Specifically, in Option 7.2a, the PHY functions belonging to O-RU are beamforming, inverse fast Fourier transform (iFFT), and cyclic prefix insertion. The rest of the PHY functions are in the O-DU. In Option 7.2b, the precoding function is added to the O-RU. As the FFT function is deployed in the O-RU, the FH data rate and required bandwidth are significantly reduced compared to Option 8. The Option 7.2x enables a relatively simple RU whose size and power consumption support network densification and enable sharing by multiple operators, which is better suited for macrocell deployment in urban areas in terrestrial networks [andersson2021functional].
@@
The different FH split option choices impact the network via three aspects, including the data rate requirement, the latency requirement, and the cooperation scale. Each of these aspects is examined in detail as follows.
@@
Data Rate: The required FH data rate increases significantly when the functional split occurs in lower PHY layers, due to the need to transmit raw In-phase and Quadrature (IQ) samples or partially processed signals [larsen2018survey]. Apart from that, the data rate is also influenced by the number of bits required to represent matrix elements in the precoding and beamforming functions, as it may vary depending on the dynamic range and precision requirements given the underlying matrix multiplication operations in lower PHY layers. Therefore, both FH split options and the quantization effect should be considered to satisfy the FH data rate constraint [park2014fronthaul].
@@
The FH data rate constraint in NTN is more stringent than in traditional TN, where the optical fiber is typically deployed between O-RU and O-DU. However, in NTNs, the functional entities (i.e., O-DU and O-RU) may be deployed on different platforms depending on the payload capabilities, making it challenging to employ wireline-based FH solutions. For example, the authors in [lee2025ran] proposed a LEO-based NTN architecture with the O-CU located on the ground, while the RU and DU were deployed on separate satellites. The FH transmission was achieved via inter-satellite links (ISLs) by optimizing signal compression. In addition, the authors in [zhang2021optimizing] considered the free-space optical (FSO) communication for the FH link due to its high data rate and bandwidth. Nevertheless, FSO-based FH remains challenged by atmospheric turbulence and the beam alignment problem, particularly for mobile NTN platforms such as HAPs or UAVs operating in more complicated atmospheric conditions. Therefore, to guarantee the stringent FH data rate constraint, it is essential to carefully design the functional split option and compression techniques tailored to the specific characteristics of different NTN platforms.

### v3

@@ 3-A NTN SPECIFICATIONS AND CONSTRAINTS
The typical weights of macro massive MIMO RU and baseband unit (BBU) are over 10Kg and 5Kg, respectively [45]. Noticeably, in the 5G era, the maximum power consumption of a 64T64R active antenna unit (AAU) is 1000–1400 W, and that of a BBU is about 2000 W [60]. However, different NTN platforms have various specifications and constraints, which should be considered in the NTN development stage [54, 112].
@@
The DJI flagship untethered UAV platform Matrice 350 RTK achieves a loading capability of 2.7 Kg with 31 minutes of flight time, which cannot carry the macro BS equipment discussed above [42]. Although DJI has customized a delivery UAV platform with a 40 Kg loading capability (i.e., DJI Flycart30). The flight time is limited to 8 minutes, which fails to guarantee continuous connections to the users [41]. In addition, the battery is originally designed to support the UAV flight with a typical power supply of 0.3 kWh (e.g., TB65 from DJI). Apart from untethered UAVs, tethered UAVs have also been regarded as an emerging NTN platform due to their tethered systems. The DG-M20 tethered UAV, from Dagong Technology Co., Ltd, can carry 15 Kg payloads hovering at 100 m for 24 hours [34]. Furthermore, the HAP has also attracted attention from both academia and industry due to its relatively lower altitude and quasi-stationary position. Taking the HAP from Stratospheric Platforms Ltd as an example, with a wingspan of 60 m, it can support a maximum of 140 Kg payload and can provide a 20 kWh power supply [1]. The typical specifications of NTN platforms are summarized in Table 3, including loading capability, power supply, and other related aspects.
@@
To overcome the above specification limitations, the standardized function splits and open interfaces defined in ORAN can be leveraged. These mechanisms provide flexibility for NTN MNOs to customize onboard components O-RU/O-DU/O-CU from different vendors, thereby meeting the specification requirements of NTN platforms. For example, the commercial untethered UAV can be utilized to implement a micro or pico network with integrated O-RU/O-DU/O-CU and 5GC on board, which acts as an E2E flexible deployment option and avoids adding peripheral equipment for FH, midhaul, and backhaul transmissions. On the contrary, the power cable and optical fiber in the tethered UAV system allow it to serve as a macro network with various onboard options, including O-RU, O-RU+O-DU, and O-RU+O-DU+O-CU, which can be configured according to the latency requirement of the service.
3-B NTN ROBUST CONNECTIVITY
@@
Due to the mobility of the NTN platforms and the varying channel conditions, one important challenge is to guarantee robust connectivity during operation, especially for the LEO satellites with high speeds. To solve this challenge, a simple solution is to deploy a massive constellation of satellites, which can provide redundant access links and transportation links under a highly dynamic network topology. For example, SpaceX from the USA has successfully launched more than 5000 Starlink satellites into orbit. In addition, China has proposed GW-A59 and GW-2 sub-constellations as GuoWang (national network) mega-constellation, a large satellite internet project aimed at providing broadband services. It is worth noting that GW-A59, with 6,080 satellites, is intended to orbit at a lower altitude (below 500 km), while GW-2, with 6,912 satellites, will orbit at a higher altitude (1,145 km above Earth). However, the massive deployment of satellites imposes a heavy financial investment burden on a single NTN MNO and requires long development periods.
@@
The integration of ORAN’s interoperability into NTN has the potential to bring new opportunities to solve the challenge above by enabling network sharing among MNOs. At the end of 2024, 3GPP identified Indirect Network Sharing (INS) as the next evolution step towards 5G-Advanced Rel-20, which aims to enhance network co-construction and sharing capability [6]. Specifically, INS was regarded as an effective solution to mitigate the compatibility issues between the network elements of sharing parties, e.g., shared satellite access network and two or more core networks of terrestrial participating operators via multiple operator core network (MOCN) [2]. However, to avoid direct connections between network elements of different MNOs, INS is based on the build of a hosting network in advance by all participating MNOs, which cannot adapt to the high mobility and complex NTN environments. ORAN surpasses the INS concept by enabling direct interconnections among MNOs, and has the potential to achieve real-time on-demand sharing based on the instantaneous network status, thereby offering greater flexibility in NTN. For clarity, consider a scenario in which O-RU, O-DU, and O-CU components of MNO A experience failures due to space weather events. In this case, MNO A can temporarily access the O-RU, O-DU, and O-CU resources of MNO B through the E2, E1, F1, and Open FH interfaces to restore service continuity. Meanwhile, to strengthen the interface security during on-demand sharing among MNOs, O-RAN Alliance has also defined dedicated security mechanisms for each interface [93]. Furthermore, the virtualization of ORAN enables the deployment of Docker technology, which can also enhance data isolation and security during data processing among MNOs.
3-C NTN PLATFORM CONTROL AND DATA ANALYTICS
@@
In traditional TNs, the BS is usually installed on fixed terrestrial infrastructure with a sufficient and stable power supply, such as cell towers, where the network performance is independent of the terrestrial infrastructure during operation. However, in the NTN scenario, the instantaneous position, velocity, and battery status of the platforms impact the network performance significantly during handover, beam management, or power allocation. To improve the network performance, it is important to design an interface between the NTN platform and onboard communication components. This interface can support dynamic optimization of communication algorithms by adapting to instantaneous platform states, and also enable proactive platform control. However, following the vendor-locked principle in traditional RAN, the vendor is required to develop both the NTN onboard components and the NTN platform to ensure compatibility, which may increase development complexity and investment costs.
@@
The recent O-RAN Alliance white paper in [39] has proposed integrating ORAN into NTNs to address these challenges by providing standardized and open interfaces between the RIC/SMO and NTN platform. However, it mainly focuses on offloading the additional terrestrial traffic, instead of designing the open interfaces between the onboard component and the NTN platform for real-time control and optimization. Therefore, further standardization of open interfaces between the NTN platform and O-RU/O-DU/O-CU is essential to enhance network performance and deployment flexibility. For example, DJI has provided standard application programming interface (API) functions to control UAV mobility and access real-time UAV status. Specifically, the Roll, Pitch, and Yaw information and battery status can be accessed periodically under pre-defined binary control and command (C&C) formats. In addition, the design of such interfaces also requires close collaboration between NTN platform manufacturing communities and the O-RAN Alliance to accurately capture the diverse characteristics of NTN platforms.
3-D NTN ON-ORBIT TEST AND UPGRADE
@@
Traditional RAN is mainly based on vendor-locked hardware, and most network functions and algorithms are implemented on specialized hardware with limited upgrade capability, such as application specific integrated circuit (ASIC), field-programmable gate array (FPGA), and digital signal processor (DSP). The specialized hardware is effective in TN due to the easy access and low upgrade costs for testing and upgrading. However, in the NTN scenario, launching NTN platforms for the testing or deployment of new features is much more costly and time-consuming [107, 76]. Therefore, it is essential to achieve on-orbit testing and upgrading to improve flexibility and shorten DevOps cycles in NTNs. In this regard, based on network function virtualization (NFV), ORAN enables the algorithms to be deployed on commercial off-the-shelf (COTS) hardware, such as central processing unit (CPU) or graphics processing unit (GPU). This virtualization capability will allow MNOs to test and upgrade the related components in NTN in a more agile DevOps manner.
@@ 4 AI-NATIVE ORAN-BASED NTN FRAMEWORK
Motivated by the advantages of ORAN-based NTN and the envisioned AI-based 6G networks, in this section, we propose an AI-Native ORAN-based NTN framework. The architecture details and key technological enablers of the proposed framework are presented below.
Figure 6: Proposed orchestrated AI-Native ORAN-based NTN framework.
4-A ARCHITECTURE
@@
The proposed framework is illustrated in Fig. 6, which spans the RAN, transport network (TPN), and core network (CN) domains. Specifically, the RAN domain is implemented within the NTN segment, while the TPN domain includes both NTN and TN segments, and the CN domain is deployed in the TN segment. Each domain is managed and orchestrated by a controller, which is equipped with AI models and digital twin technology to enhance data analysis and predictive capabilities.
The cross-domain SMO includes three domain controllers and a cross-domain AI controller. It is also responsible for O-Cloud management and orchestration, and RAN network function operation and maintenance as defined by the O-RAN Alliance. The domain controllers, including Non-RT RIC, TPN domain controller, and CN domain controller, host AI models to infer the optimized variables in RAN, TPN, and CN domains, respectively. Each domain controller can access the domain-specific data for AI model training and execution. Due to the hierarchical timescale characteristic, the TPN domain controller can access historical data from the RAN domain, and the CN domain controller can further utilize historical data from both the RAN and TPN domains.
Beyond single-domain intelligence, a cross-domain AI controller is introduced at the cross-domain SMO layer for AI lifecycle management. This controller can operate in two modes. For the first mode, the cross-domain AI controller serves as a central server, which helps facilitate the coordination of MADRL or aggregate the local models for a global model in FL. For the second mode, the cross-domain AI controller aims to achieve E2E optimization, and it hosts an AI model to infer the optimal policy based on the three single-domain controllers, thereby guaranteeing service-level agreements (SLAs).
It is worth noting that the proposed framework is designed to be an AI-Native system by supporting each AI-Native aspect defined in Section 2-C, and the details of the realization of each AI-Native aspect are presented in Table 4. Furthermore, the proposed framework consists of four key technological enablers, including dynamic wireless FH split, multi-tier RIC structure, flexible functional role switch, and cross-domain SMO. Each key enabler helps facilitate the implementation and operation of this proposed framework, which will be discussed in detail in the following subsections.
Table 4: Detailed realization of AI-Native aspects in the proposed ORAN-based NTN Framework
AI-Native Aspect Realization in the Proposed ORAN-based NTN Framework
Intelligence Everywhere Multi-tier deployment of AI models across Non-RT RIC, Near-RT RIC, and RT RIC to support decision-making at different timescales. Cross-domain AI controller supports coordinated intelligence across RAN, TPN, and CN domains.
Distributed Data Infrastructure Supports cross-layer and cross-domain data and knowledge sharing through open interfaces of ORAN and cross-domain SMO. Provides distributed computing resources by enabling edge computing on NTN platforms and cloud computing in the cross-domain SMO or Non-RT RIC.
Zero-touch Automation AI-assisted FCAPS operation in cross-domain SMO enables the self-configuration, self-healing, self-optimization, and self-protection. xApps and rApps form closed-loop control mechanisms for near-real-time and non-real-time self-optimization. The dynamic wireless FH split and flexible functional role switch enables the self-configuration to adapt to rapidly changing NTN scenarios.
AI as a Service (AIaaS) Cross-domain AI controller enables shared access to AI-based inference services across RAN, TPN, and CN. The AI lifecycle management and computing resource of this framework can be utilized by third parties via open interfaces for their AI model training or execution.
4-B ENABLER-1: DYNAMIC WIRELESS FH SPLIT
@@
The gNB’s operation is conceptually structured as a series of functions, with the distribution of these functions defined by functional splits. Within the 3GPP framework, eight distinct functional split options exist, as shown in Fig. 7. Each split option describes how the logical nodes interrelate to one another and what specific activities each undertakes [10]. From Option 1 to 8, there is a higher degree of centralization and a notable improvement in resource management efficiency, which leads to reduced O-RU complexity and associated costs. However, this comes at the expense of imposing high data rates and strict latency requirements on the FH between O-DU and O-RU [70].
@@
Typically, the O-RAN Alliance lends its support to the Option 7.2x variant for networks with high-capacity and high-reliability requirements, and this variant further splits the functions in the L-PHY based on the PHY split in Option 7. This option is made possible by using the eCPRI interface for FH. In Option 7.2x, the O-CU/O-DU manages RRC, PDCP, RLC, MAC, H-PHY, and parts of L-PHY, while the O-RU handles RF and the rest of L-PHY. The ’x’ in this split name can be either ’a’ or ’b’, depending on whether some of the PHY functionalities reside in H-PHY or L-PHY [128]. Without loss of generality, our following discussion is based on the downlink transmission function split. As shown in Fig. 7, for Option 7.2a, the PHY functions belonging to O-RU are beamforming, inverse fast Fourier transform (iFFT), and cyclic prefix insertion. The rest of the PHY functions are in the O-DU. In Option 7.2b, the precoding function is added to the O-RU. As the FFT function is deployed in the O-RU, the FH data rate and required bandwidth are significantly reduced compared to Option 8. The Option 7.2x enables a relatively simple RU whose size and power consumption support network densification and enable sharing by multiple operators, which is better suited for macrocell deployment in urban areas in terrestrial networks [19].
@@
The different FH split option choices impact the network via three aspects, including the data rate requirement, the latency requirement, and the cooperation scale. Each aspect is examined in detail as follows.
4-B.1 Data Rate
@@
The required FH data rate increases significantly when the functional split occurs in lower PHY layers, due to the need to transmit raw In-phase and Quadrature (IQ) samples or partially processed signals [70]. In addition, the data rate is also influenced by the number of bits required to represent matrix elements in the precoding and beamforming functions, as it may vary depending on the dynamic range and precision requirements given the underlying matrix multiplication operations in lower PHY layers. Therefore, both FH split options and the quantization effect should be considered to satisfy the FH data rate constraint [96].
@@
The FH data rate constraint in NTN is more stringent than in traditional TN, where the optical fiber is typically deployed between O-RU and O-DU. However, in NTNs, O-RU and O-DU may be deployed on different platforms depending on the payload capabilities, making it challenging to employ wireline-based FH solutions. For example, the authors in [71] proposed a LEO-based NTN architecture with the O-CU located on the ground, while the RU and DU were deployed on separate satellites. The FH transmission was achieved via inter-satellite links (ISLs) by optimizing signal compression. In addition, the authors in [134] considered the free-space optical (FSO) communication for the FH link due to its high data rate and bandwidth. Nevertheless, FSO-based FH remains challenged by atmospheric turbulence and the beam alignment problem, particularly for mobile NTN platforms such as HAPs or UAVs operating in more complicated atmospheric conditions. Therefore, to guarantee the stringent FH data rate constraint, it is essential to carefully design the functional split option and compression techniques tailored to the specific characteristics of different NTN platforms.